In the reaction N2+3H2>2NH3, what is the mole ratio of hydrogen to nitrogen


1
Expert's answer
2022-05-27T14:45:03-0400

N2 + 3H2 = 2NH3

The ratio of the coefficients in front of H2 and N2 corresponds to their ratio. that is, 3 and 1

Mole ratio of hydrogen to nitrogen =3 / 1 =3

ANSWER: 3